OpenStudy (omgthatgirl):

HELP Solve the equation 16x^2+9=0 a. 3/4 b. 3/4i c. 4/3i d. 9/16i

OpenStudy (amorfide):

rearrange to make the equation equal to a constant \[16x^{2}+9=0\] \[16x^{2}=-9\] solve for x \[x^{2}=\frac{ -9 }{ 16 }\] \[x=\sqrt{(\frac{ -9 }{ 16 })} \] \[\sqrt{\frac{ a }{ b }}=\frac{ \sqrt{a} }{ \sqrt{b} }\] \[x=\frac{ \sqrt{-9} }{ \sqrt{16} }\] \[\sqrt{-a}=\sqrt{-1}\sqrt{a}\] \[\sqrt{-1}= i\] \[x=\frac{ \sqrt{-1}\sqrt{9} }{ \sqrt{16} }\]

OpenStudy (omgthatgirl):

so its d? @amorfide

OpenStudy (amorfide):

no I didn't finish it, I left it for you to finish you know sqrt9 is 3 sqrt -1 is i sqrt 16 is 4 simplify what I gave you and you get your answer

OpenStudy (omgthatgirl):

3/4i

OpenStudy (amorfide):

well done

OpenStudy (omgthatgirl):

thanks :)